Q.18 A vessel has 4l and 500ml of curd. In how many glasses each of 25ml capacity, can it be filled?

Grade:6

1 Answers

Armaan Arora
2 years ago
4l 500ml = (4*1000 +500)ml =4500ml
capacity of each glass = 25ml 
no of glasses that can be filled= 4500/25 =180 
so 180 glasses can be easily filled out of it